OpenStudy (anonymous):

what is the simplified form of 15x^8 4x^4 ------ divided by ------ ? 24y^5 8y^2

OpenStudy (turingtest):

\[\frac{15x^8}{24y^5}\div\frac{4x^4}{8y^2}\]

OpenStudy (turingtest):

this?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

use\[\frac ab\div\frac cd=\frac ab\times\frac dc=\frac{ad}{bc}\]

OpenStudy (turingtest):

(i.e. division is the same as multiplication by it's reciprocal)

OpenStudy (turingtest):

caveat: not true for \(b,d,c=0\)

OpenStudy (anonymous):

How would you simplify this, I'm stuck

OpenStudy (turingtest):

try doing what I wrote: flip the second fraction and make the "divided by" into "times"

OpenStudy (anonymous):

I'm ok with that I'm stuck on the part where you factor

OpenStudy (turingtest):

write what you have so far so I know what you mean use parentheses like (15x^2/24y^5)/(4x^4/8y^2)=...

OpenStudy (anonymous):

(15x^8/24y^5)(4x^4/8y^2), then keep-change-flip: (15x^8/24y^5)(8y^2/4x^4), then factor

OpenStudy (turingtest):

since you are multiplying the numerators and denominators are laterally interchangeable\[\frac{ab}{cd}=\frac{ba}{cd}\]so you can write (15x^8/24y^5)(8y^2/4x^4) =(15x^8/4x^4)(8y^2/24y^5) now you can use the rule\[\large\frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (anonymous):